# MAESTRO: Multi-Agent Environment Shaping through Task and Reward Optimization

### v1

@@ 4.2 LLM-Driven Adaptive Curriculum
The adaptive curriculum maintains a scalar difficulty $d\in[0.3,1.0]$ and updates it based on a rolling window of five episodes using the current policy’s cumulative return. Difficulty adjustment follows:
  • If the average return exceeds $1.05\times 165.0$ (173.25) for two consecutive windows, increase $d$ by $5\%$.
  • If the return drops below $0.95\times 165.0$ (156.75), decrease $d$ by $10\%$.
  • If $d$ remains unchanged for 15 episodes, apply a forced $5\%$ increase (plateau escape).
Difficulty is clipped to $[0.3,1.0]$ to avoid trivial (empty-traffic) or over-saturated scenarios.
Under llm_adaptive (A2, A7), GPT-4o-mini is prompted with the current difficulty $d$ and summary statistics to generate a semantically rich traffic context (e.g., flow patterns, demand asymmetry, vehicle parameters), which is then converted into a CityFlow configuration. Under llm (A8), a less constrained prompting and regeneration strategy is used for comparison, while applying the same difficulty-update rules. Pilot studies verified that increasing $d$ results in monotonic increases in fundamental difficulty metrics (queue length, delay) under a fixed-time policy.
4.3 LLM-Generated Reward Shaping (MAESTRO – A7)
The full MAESTRO configuration (A7) extends the A2 curriculum baseline by introducing LLM-generated auxiliary rewards that evolve with the curriculum.
Reward Generation and Validation.
At regeneration events (triggered by $\Delta d\geq 0.15$, 20 episodes since the last generation, or a stagnation-induced difficulty increase), GPT-4o-mini is prompted with the current difficulty, aggregated statistics, and a description of the learning goal. The LLM generates a short Python reward function $r_{\mathrm{LLM},t}$. A three-stage validation pipeline ensures safety:
  • Syntax Check: abstract syntax tree validation,
  • Sandboxed Execution: sample rollouts to confirm runtime behavior,
  • Safety Check: range bounding and finite-value verification.
If three consecutive generations fail, the system reverts to environment-only rewards (behaving as A2) until the next regeneration trigger.
Scaling and Difficulty-Dependent Weighting.
Because LLM-generated rewards can differ in scale, we apply a fixed scaling factor of $170.0$ to match typical environment reward magnitudes. A difficulty-dependent weight
$w(d)=w_{\min}+\frac{d-d_{\min}}{d_{\max}-d_{\min}}(w_{\max}-w_{\min}),$
with $w_{\min}=0.1$, $w_{\max}=0.5$, $d_{\min}=0.3$, and $d_{\max}=1.0$, controls the auxiliary contribution. This ensures low influence at easy difficulties and stronger guidance in challenging regimes.
Combined Reward.
The per-timestep reward used for MADDPG updates is:
$r_{t}=\begin{cases}r_{\mathrm{env},t},&\text{A2, A8 (Curriculum Baselines)},\\[3.0pt] (1-w(d))\,r_{\mathrm{env},t}+w(d)\cdot 170.0\cdot r_{\mathrm{LLM},t},&\text{A7 (Full MAESTRO)}.\end{cases}$
Here, $r_{\mathrm{env},t}$ is the CityFlow reward and $r_{\mathrm{LLM},t}$ is the auxiliary shaping signal generated by the Architect.
5 Experimental Design and Evaluation Protocol
We evaluate the MAESTRO framework by quantifying the contribution of LLM-generated reward shaping and adaptive curriculum design to the performance and robustness of cooperative MARL. Our goals are twofold: (1) isolate the effect of LLM-generated reward functions relative to template-based rewards under an adaptive curriculum, and (2) demonstrate the stability and learning efficiency of the full MAESTRO system in a complex, non-stationary environment.
Experimental Conditions.
We evaluate three configurations forming a controlled ablation (detailed in §4):
  • A2 (Curriculum Baseline): Uses the llm_adaptive curriculum with a fixed template-based environment reward.
  • A8 (Alternative Curriculum): Uses an alternative llm curriculum implementation and the same template-based reward.
  • A7 (Full MAESTRO): Uses the A2 curriculum but replaces template rewards with dynamic LLM-generated reward functions, scaled by 170.0 and weighted by a difficulty-adaptive factor $w\in[0.1,0.5]$.
All experiments use identical hyperparameters, network architecture, agent design, and simulation settings.
Key Comparisons.
This design supports two central comparisons: (1) A2 vs. A7 isolates the contribution of LLM-generated reward shaping within the same curriculum framework; (2) A2 vs. A8 evaluates two LLM curriculum strategies while keeping the reward design fixed. We train all three configurations for 200 episodes using four seeds, prioritizing deeper runs to strengthen statistical power and attribute performance differences to specific LLM components.
5.1 Environment and Curriculum Setting
All experiments use the Hangzhou (HZ) traffic network with 16 signalized intersections (§3.2). Each episode spans 360 steps. We report standard traffic-signal-control metrics: mean delay, queue length, throughput, and cumulative return.
The adaptive curriculum maintains a scalar difficulty $d$ updated from a rolling five-episode window. When the windowed return exceeds 173.25 (105% of a target 165.0) for two consecutive windows, $d$ increases by 5%. When it falls below 156.75 (95% of target), $d$ decreases by 10%. If unchanged for 15 episodes, $d$ increases by 5% to escape stagnation. The difficulty is clipped to $[0.3,1.0]$. Difficulty monotonicity with respect to task hardness (queue length and delay) is empirically verified under a fixed-time control policy. Because difficulty updates are performance-driven, curriculum trajectories differ by seed.
@@ 5.1 Environment and Curriculum Setting
In the full MAESTRO system (A7), difficulty changes and stagnation events also trigger LLM reward regeneration (§4.3), whereas A2 and A8 apply identical curriculum logic but omit reward regeneration.
@@ 5.2 Metrics
We evaluate results along three dimensions:
  • Traffic Efficiency and Final Performance: Episode return, mean delay, queue length, and throughput measured over episodes 180–199.
  • Learning Efficiency and Stability: Episodes required to reach the target performance (165.0), time to reach 90% of final return, and stability measured by variance and standard deviation of returns in episodes 180–199, supplemented with a risk-adjusted metric (Sharpe Ratio).
  • Curriculum Behavior: Per-seed difficulty trajectories, counts of difficulty adjustments, and final difficulty reached.
5.3 Experimental Protocol
Each configuration (A2, A8, A7) is trained with four seeds (200, 300, 400, 500). Training runs for 200 episodes (72,000 environment steps). Evaluation episodes occur every two training episodes, with deterministic policies (no exploration noise) to measure true policy quality. Hyperparameters follow standard MADDPG practice, and all simulation parameters (network topology, demand, and CityFlow settings) remain fixed.
Reproducibility is enforced by seeding PyTorch, NumPy, Python RNG, and CuDNN. In A7, LLM-generated reward code is processed through a deterministic validation and post-processing pipeline, ensuring that identical seeds lead to identical code acceptance decisions.
Statistical analyses operate on seed-level aggregates. For each configuration and seed, we compute mean return over episodes 180–199 as the summary of final performance. Pairwise comparisons use the Mann–Whitney U test ($\alpha=0.05$, Bonferroni corrected). Effect sizes use Cohen’s $d$, and uncertainty is expressed via 95% bootstrap confidence intervals computed from 10,000 resamples. We avoid pooling episodes across seeds to preserve independence and capture between-seed variance.
5.4 Visualization Outputs
Results are visualized through four key figures: (1) learning curves showing episode return over training with 95% confidence bands (Fig. 3); (2) box plots with per-seed markers summarizing final performance (Fig. 4); (3) ablation study comparing component contributions across A2, A8, and A7 (Fig. 5); and (4) stability analysis via coefficient of variation comparison (Fig. 6).
@@ 6 Results
We evaluate MAESTRO through a comprehensive analysis of learning dynamics, final performance, traffic efficiency, and training stability across the three configurations: A2 (Curriculum Baseline), A7 (Full MAESTRO), and A8 (Alternative Curriculum). Figure 3 illustrates the learning curves, Figure 4 presents final performance distributions, and Table 2 summarizes the quantitative outcomes.
6.1 Traffic Efficiency Improvements
Both MAESTRO-enhanced configurations (A7, A8) outperform the baseline A2 in late-training traffic efficiency metrics. Table 1 reports representative results for seed 200. A8 yields the largest improvements, reducing mean delay by 27.8% and increasing throughput by 9.2%. A7 also demonstrates substantial gains, decreasing delay by 19.9%. These improvements confirm that both curriculum modulation (A8) and reward shaping (A7) provide meaningful benefits in traffic throughput and congestion mitigation.
Table 1: Traffic efficiency metrics during late-training episodes (seed 200). Higher throughput and lower temporal metrics indicate better performance. Percentages denote relative changes compared to A2.
@@ 6.1 Traffic Efficiency Improvements
Figure 3: Training performance across 200 episodes. Solid lines denote mean return; shaded regions indicate 95% confidence intervals. A7 shows the narrowest confidence band, indicating the highest training stability.
6.2 Final Performance and Stability Analysis
Table 2 reports final performance statistics (episodes 180–199). A8 achieves the highest mean return (160.35), but with markedly higher variance (CV 2.55%), consistent with the wide spread seen in Figure 4. In contrast, A7 attains moderate performance gains (158.87) with exceptionally low variance (CV 0.65%), reflecting highly consistent learning dynamics across seeds.
Table 2: Ablation summary of episode returns over episodes 180–199 (higher is better). CV: coefficient of variation. Sharpe Ratio (SR) = (mean – A2 mean)/std.
@@ 6.2 Final Performance and Stability Analysis
A2 (Baseline) llm_adaptive Template $157.04\pm 2.66$ 1.69 159.63 152.82
A7 (Full MAESTRO) llm_adaptive LLM-shaped $158.87\pm 1.04$ 0.65 1.76 160.15 157.76
A8 (Alt. Curric.) llm Template $160.35\pm 4.08$ 2.55 0.81 166.05 154.77
A key observation is that MAESTRO’s reward shaping mechanism (A7) yields robust, risk-adjusted improvements. While A8 provides stronger peak performance, its variability makes it less reliable. A7 achieves a Sharpe Ratio of 1.76 (more than twice A8’s 0.81), driven by consistent performance across all seeds.
Figure 4: Distribution of final episode returns (episodes 180–199). A7 exhibits tight clustering around the mean, whereas A8 shows large variability, consistent with Table 2.
@@ 6.3 Ablation Insights
Switching from llm_adaptive (A2) to llm (A8) yields a notable increase in mean return, confirming that more expressive curriculum descriptions can produce harder and more informative training samples. However, the substantially higher variance indicates that unconstrained curriculum generation can produce unstable learning trajectories.
@@
Introducing LLM-generated reward functions yields consistent gains under an identical curriculum schedule. The dramatically reduced variance of A7 supports the interpretation that learned reward functions act as structured priors, damping out high-variance exploration and stabilizing policy updates.
@@
A7 and A8 illustrate a fundamental trade-off: A8 delivers higher peak performance but at the cost of instability, while A7 provides smaller but highly reliable improvements. For safety-critical or real-world deployment, the robustness of A7 is preferable.
Figure 5: Mean returns and standard deviations across configurations. A8 offers the largest absolute gain (+3.31 over A2), whereas A7 achieves smaller gains with significantly improved stability.
@@ 7.1 Key Findings and Methodological Implications
Curriculum Design as the Primary Driver.
The most pronounced performance gains arise from modifying the curriculum-generation strategy (A2 $\rightarrow$ A8). This highlights that task selection and progression remain the dominant levers in determining MARL efficiency and asymptotic performance. The ability of the LLM to synthesize semantically rich and diverse scenarios enables more informative training distributions than manually designed or template-based curricula.
Modest Return Gains, Large Domain Impact.
Although aggregate episode-return improvements appear modest (1–2%), domain-specific metrics reveal large operational benefits: 20–28% reductions in delay and 6–9% improvements in throughput. These findings reinforce the importance of evaluating MARL systems not only through abstract learning metrics but also through domain-relevant indicators that capture real-world utility.
7.2 Limitations and Future Work
Generalization Beyond Traffic Control.
Experiments are conducted exclusively in the Hangzhou TSC environment. Future work should investigate MAESTRO in broader cooperative MARL domains—such as multi-robot coordination, distributed resource allocation, and networked systems—to assess whether the LLM-as-Architect paradigm generalizes to structurally different environments.
Dependence on Model and Prompt Design.
The framework relies on a specific model (GPT-4o-mini) and a fixed prompt configuration. Systematic prompt optimization, prompt diversification, or incorporating few-shot demonstrations may yield higher-quality generative outputs. Evaluating performance across different LLM families would help clarify the robustness of MAESTRO to model choice.
Richer Curriculum and Reward Structures.
An important direction is expanding the expressiveness of the curriculum space and reward generator. Integrating LLM-informed priors into policy or value initialization, incorporating temporal abstractions, or conditioning curriculum generation on uncertainty estimates could further enhance sample efficiency and generalization.
Safety and Interpretability of Generated Rewards.
Although MAESTRO employs syntactic, runtime, and safety checks for generated reward code, subtle misalignment or specification errors remain possible. Future work should explore formal verification of LLM-generated code, human-in-the-loop oversight, and runtime safety filters to ensure alignment with operational and ethical constraints, especially for real-world deployment scenarios.

### v2

@@ 4.2 Adaptive Curriculum
The curriculum maintains a scalar difficulty $d\in[0.3,1.0]$, updated from a rolling window of five episodes using cumulative return. Let $\bar{R}_{w}$ denote the windowed mean and $\tau=165.0$ a fixed threshold. The update rule is:
[…]
The asymmetric patience (2 for increases, 1 for decreases) favors stability by requiring sustained success before progression while quickly reacting to deterioration. Difficulty bounds avoid trivial or saturated regimes.
Under llm_adaptive, GPT-4o-mini receives the current $d$ and summary statistics and outputs a natural-language traffic context; this is compiled into CityFlow config files. Under llm, prompts are less constrained but share the same difficulty updates. Pilot experiments under a fixed-time controller confirm that increasing $d$ monotonically increases difficulty (queue length, delay).
The threshold $\tau=165.0$ corresponds to the mean return of a competent fixed-time baseline over 200 episodes and is stable (variance <<3%). Multipliers $1.05$ and $0.95$ create hysteresis around this anchor.
Difficulty is mapped to CityFlow parameters via linear interpolation; parameters that increase difficulty (e.g., maximum speed, acceleration) grow with $d$, while others (e.g., minimum gap, headway) shrink. Each parameter is further perturbed by $\pm 20\%$ multiplicative noise to diversify scenarios at fixed difficulty.
4.3 Training Algorithm
Algorithm 1 summarizes MAESTRO. The outer loop samples curriculum tasks and updates difficulty; the inner loop performs MADDPG updates using either environment-only rewards (A2/A8) or combined environment + LLM shaping (A7).
Algorithm 1 MAESTRO: Dual-Loop Training with Adaptive Curriculum and LLM Guidance
1:Initialize difficulty $d\leftarrow 0.3$, window $W\leftarrow\emptyset$, policy $\pi_{\theta}$, replay buffer $\mathcal{D}$
2:Initialize LLM reward $r_{\text{LLM}}\leftarrow\text{None}$, prior policy $\pi_{\text{LLM}}\leftarrow\text{None}$
3:Set threshold $\tau=165.0$, patience $(p^{+},p^{-})=(2,1)$, regularization schedule $\alpha(t)$
4:while $t_{\text{env}}<T_{\text{max}}$ do $\triangleright$ Outer loop
5:  task $\leftarrow\text{GenerateTask}(d)$ $\triangleright$ LLM-generated context
6:  Reset environment with task; $\tau_{\text{episode}}\leftarrow\emptyset$
7:  for $t=1$ to $T_{\text{episode}}$ do
8:    $a_{t}\sim\pi_{\theta}(s_{t})$ with exploration noise
9:    $r_{\text{env},t},s_{t+1}\leftarrow\text{Environment.step}(a_{t})$
10:    if $r_{\text{LLM}}\neq\text{None}$ then $\triangleright$ A7 only
11:     $w\leftarrow 0.1+0.4\cdot d$
12:     $r_{t}\leftarrow r_{\text{env},t}+w\cdot r_{\text{LLM},t}(s_{t},a_{t})$
13:    else
14:     $r_{t}\leftarrow r_{\text{env},t}$
15:    end if
16:    Append $(s_{t},a_{t},r_{t},s_{t+1})$ to $\tau_{\text{episode}}$
17:  end for
18:  Store $\tau_{\text{episode}}$ in $\mathcal{D}$
19:  $R_{\text{episode}}\leftarrow\sum_{t}r_{\text{env},t}$; update window $W$
20:  if $|W|\geq 5$ then $\triangleright$ Difficulty update
21:    $\bar{R}\leftarrow\text{mean}(W)$
22:    Update $d$ using thresholds $1.05\tau$ and $0.95\tau$ with hysteresis and anti-stagnation
23:  end if
24:  if A7 and regeneration trigger fired then
25:    $(r_{\text{LLM}},\pi_{\text{LLM}})\leftarrow\text{LLM-Generate}(d,W,\text{task})$
26:    Validate templates via syntax, sandbox, and safety checks
27:  end if
28:  if $|\mathcal{D}|\geq B_{\text{batch}}$ then $\triangleright$ Inner loop
29:    for $k=1$ to $K_{\text{updates}}$ do
30:     Sample minibatch $\mathcal{B}$; update critic by TD error
31:     Compute actor loss $\mathbb{E}[Q_{\phi}(s,\pi_{\theta}(s))]$
32:     if A7 then
33:      Add MSE regularization $\alpha(t)\|\pi_{\theta}(s)-\pi_{\text{LLM}}(s)\|_{2}^{2}$
34:     end if
35:     Update $\theta,\phi$ and their targets via soft updates
36:    end for
37:  end if
38:end while
4.4 LLM-Generated Reward Shaping (A7)
A7 extends A2 by adding LLM-shaped reward and prior policy regularization.
Template-based reward generation.
At regeneration events (difficulty change $\Delta d\geq 0.15$, 10 episodes since last generation, or template switch, with a minimum of 3 episodes between generations), GPT-4o-mini is prompted with the current difficulty, performance statistics, and high-level goals. Instead of emitting raw Python code, the LLM outputs parameters for pre-validated templates (e.g., weights over delay, queue length, phase switching). The templates enforce weight normalization, bound outputs, and prevent NaNs/Infs. A three-stage validation pipeline (syntax, sandboxed execution, safety checks) filters out invalid generations. If three consecutive generations fail, the system falls back to environment-only rewards until the next trigger.
Prior policy regularization.
The LLM also parameterizes a prior policy $\pi_{\text{LLM}}(a|s)$, represented as logits over actions. This prior is never used for action selection or initialization; the actor $\pi_{\theta}$ always chooses actions and starts from random weights. The prior is first generated after an exploration episode and then used to regularize the actor via:
$\mathcal{L}_{\text{reg}}(s)=\alpha(t)\,\|\pi_{\theta}(s)-\pi_{\text{LLM}}(s)\|_{2}^{2},$
with $\alpha(t)$ linearly decaying from $0.5$ at episode 1 to 0 at episode 200. The full actor loss is:
$\mathcal{L}_{\text{actor}}=-\mathbb{E}_{s\sim\mathcal{B}}[Q_{\phi}(s,\pi_{\theta}(s))]+\mathbb{E}_{s\sim\mathcal{B}}[\mathcal{L}_{\text{reg}}(s)].$
This encourages early exploration along semantically meaningful behaviors (e.g., maintaining green waves, prioritizing high-pressure links) while gradually handing over control to the learned policy.
Difficulty-dependent weighting.
The auxiliary reward weight is:
$w(d)=0.1+0.4d,$
where $d$ is a normalized difficulty score from a dedicated analyzer. Given the observed range $d\in[0.20,0.69]$, $w(d)\in[0.18,0.38]$, ensuring modest influence at easy levels and stronger shaping in harder regimes.
The combined reward is:
$r_{t}=\begin{cases}r_{\mathrm{env},t},&\text{A8},\\[3.0pt] (1-w(d))\,r_{\mathrm{env},t}+w(d)\,r_{\mathrm{LLM},t},&\text{A2, A7}.\end{cases}$
5 Experimental Design and Evaluation
We assess MAESTRO by (i) isolating the contribution of LLM-shaped rewards relative to an adaptive curriculum with template rewards, and (ii) evaluating stability and efficiency in a non-stationary TSC environment.
Conditions.
We train:
  • A2 (Curriculum Baseline): llm_adaptive curriculum, template reward.
  • A8 (Alternative Curriculum): llm curriculum, template reward.
  • A7 (Full MAESTRO): llm_adaptive curriculum, LLM-shaped rewards and prior regularization.
All use identical MADDPG architectures and hyperparameters.
Comparisons.
A2 vs. A7 isolates reward shaping; A2 vs. A8 compares curriculum modes. Each configuration is trained for 200 episodes with four seeds.
5.1 Environment and Curriculum Settings
We use the HZ network with 16 intersections; each episode has 360 steps. We report episode return and standard TSC metrics. The difficulty update mechanism is as in Section 4.2, with difficulty clipped to $[0.3,1.0]$. Difficulty trajectories differ by seed due to performance-based updates.
@@ 5.1 Environment and Curriculum Settings
In A7, difficulty changes and stagnation events also trigger LLM reward regeneration; A2 and A8 share the same curriculum logic but never generate rewards.
@@ 5.2 Metrics
We evaluate:
  • Traffic efficiency: mean delay, queue length, throughput, and cumulative return (episodes 180–199).
  • Learning dynamics: episodes to reach target performance (165.0), time to reach 90% of final return, variance and coefficient of variation (CV) of late-training returns, and a Sharpe-like risk-adjusted metric.
  • Curriculum behavior: difficulty trajectories, counts of adjustments, and final difficulty levels.
5.3 Protocol
We run each configuration with seeds 200, 300, 400, and 500. Training consists of 200 episodes (72k steps). Every second episode we run a deterministic evaluation. Random seeds control PyTorch, NumPy, Python RNG, and CuDNN. In A7, LLM template acceptance is deterministic given the seed.
For each configuration and seed, we summarize final performance by the mean return over episodes 180–199. Pairwise comparisons use Mann–Whitney U tests with Bonferroni correction; effect sizes use Cohen’s $d$, and uncertainties use 95% bootstrap confidence intervals over seeds.
5.4 Visualization
We report four plots: (1) learning curves with 95% confidence bands (Fig. 3); (2) a radar chart over normalized metrics (Fig. 4); (3) ablation bar plots for mean and variance (Fig. 5); and (4) a CV comparison (Fig. 6).
@@ 6 Results
Figure 3 shows training curves; Table 3 summarizes late-training returns.
6.1 Traffic Efficiency
Both LLM-enhanced configurations (A7, A8) improve traffic metrics over A2. Table 1 reports seed 200 as a representative example. A8 yields the largest improvements, reducing mean delay by 27.8% and increasing throughput by 9.2%; A7 also significantly reduces delay (19.9%) and wait time.
Table 1: Late-training traffic metrics (seed 200). Higher throughput and lower times are better. Percentages are relative to A2.
@@ 6.1 Traffic Efficiency
Table 2: Traffic efficiency metrics averaged across all four seeds (200, 300, 400, 500) during late-training episodes (180–199). Values reported as mean $\pm$ standard deviation across seeds. Higher throughput and lower temporal metrics indicate better performance. Percentages denote relative changes compared to A2 baseline.
Condition Throughput Avg Travel Avg Delay Avg Wait
(veh/ep) Time (s) (s) Time (s)
A2 (LLM Curr. + Template) $1570\pm 212$ $1030\pm 131$ $711\pm 143$ $763\pm 127$
A7 (Full MAESTRO) $1776\pm 188$ (+13.1%) $919\pm 116$ ($10.8\%$) $571\pm 141$ ($19.7\%$) $640\pm 127$ ($16.2\%$)
A8 (Alt. Curr. + Template) $1745\pm 302$ (+11.1%) $954\pm 169$ ($7.4\%$) $602\pm 203$ ($15.4\%$) $652\pm 215$ ($14.5\%$)
Figure 3: Training performance over 200 episodes. Lines show mean episode return; bands show 95% confidence intervals. A7 exhibits the narrowest band, indicating high stability across seeds.
6.2 Final Performance and Stability
We use episode return (cumulative reward) averaged over episodes 180–199 as the primary metric. Table 3 presents summary statistics.
Table 3: Late-training episode returns (higher is better). CV: coefficient of variation. Sharpe Ratio (SR) = (mean – A2 mean)/std.
@@ 6.2 Final Performance and Stability
A2 (Baseline) llm_adaptive Template $157.04\pm 3.07$ 1.95 159.63 152.82
A7 (Full MAESTRO) llm_adaptive LLM-shaped $158.87\pm 1.20$ 0.76 1.53 160.15 157.76
A8 (Alt. curric.) llm Template $160.23\pm 4.73$ 2.95 0.67 166.08 154.72
A8 attains the highest mean and best seed but also the highest variance. A7 achieves moderate performance gains over A2 with dramatically lower variance. Its worst seed exceeds the A2 mean, making it strongly preferable when reliability matters. Mann–Whitney U tests are underpowered at $n=4$ seeds but effect sizes are large; we treat these results as practical rather than formal significance.
Figure 4: Normalized comparison across five metrics (0–1 scale). A7 balances final return, stability, learning speed, sample efficiency, and robustness. A8 dominates in peak return but is less stable; A2 is consistently weaker.
@@ 6.3 Ablation Insights
Switching from llm_adaptive to llm increases mean return but also variance. More expressive curricula clearly help exploration but can destabilize training.
@@
Introducing LLM-shaped reward and prior regularization under the same curriculum yields consistent gains and a large reduction in variance. This supports the view that reward templates act as structured priors, smoothing exploration and updates.
@@
A7 and A8 trace a classic trade-off: A8 is best for peak performance, A7 for risk-adjusted performance. For real deployments, A7’s reliability and lower CV are more attractive.
Figure 5: Mean returns and standard deviations. A8 yields the largest raw gain; A7 offers smaller gains with much tighter error bars.
@@ 7.1 Key Findings
Curriculum is the main performance lever.
The largest jump in mean return arises from changing the curriculum mode (A2 $\rightarrow$ A8), underscoring that task selection and sequencing dominate asymptotic performance. LLM-generated semantically rich contexts make curricula a high-impact design dimension.
Small return gains, large domain impact.
While episode return improvements appear modest (1–2%), they correspond to substantial domain gains: 20–28% reductions in delay and 6–9% increases in throughput. This highlights the importance of reporting task-level metrics, not just abstract returns.
7.2 Limitations and Directions
Domain generality.
Our evaluation is limited to the HZ TSC network. Extending MAESTRO to other cooperative MARL domains (multi-robot coordination, resource allocation, network routing) is necessary to establish generality.
Model and prompt dependence.
We use GPT-4o-mini with a single prompt design. Future work should explore prompt tuning, alternative LLMs, and few-shot examples to further improve curriculum and reward quality.
Richer training signals.
Currently, MAESTRO shapes rewards and regularizes policies; future extensions could include LLM-informed value initialization, hierarchical options, or dynamic architecture adaptation.
Reward safety and interpretability.
Although template constraints and validation provide strong safeguards, subtle misalignment remains possible. For deployment, human-in-the-loop review, formal verification, or runtime monitoring will be important.